Select any two integers between -12 & +12, which will become solutions to a system of two equations. Write two equations that have two integers as solutions. Solve your system of equations, using the addition and subtraction method, showing all 5 steps. I chose 5, and -10 as my integers, so
x=5, and y=-10. My equations were:
2x + y = 0
-x + 2y =
-5 + -20 = -25
2x + y = 0
-2x + 4y =
-10 +-40 = -50
5y = -50
y = -10
Then to solve for x,
2x + y = 0
2(5)+ -10 = 0
2x =10
x= 5
Have I done this right, and included all the necessary steps as required?
Please help and show me. Thank you

Answer by Edwin McCravy(20059)

You did it correctly, but I think it could be a little clearer.  And be sure
to state your system of equations, and state exactly what you did in each step.
Here is how I would write it:

I chose 5, and -10 as my integers, so
   x=5, and y=-10.  My system of equations is:

   2x +  y = ?
   -x + 2y = ?

I find what the numbers on the right are
by substituting

   2(5) + (-10) = 10 - 10 = 0
   -(5) + 2(-10) = -5 + (-20) = -25

So my system of equations is

   2x +  y =   0
   -x + 2y = -25

To solve this system, I multiply
the second equation by 2 to
eliminate x

   2x +  y =   0
  -2x + 4y = -50
 ----------------
        5y = -50
         y = -10 

To find x, I substitute -10 for y
in the first equation and get:

   2x + (-10) = 0
      2x - 10 = 0
           2x = 10
            x = 5

Edwin